OpenStudy (anonymous):

The function fails to be differentiable at x=0. Tell whether the problem is a cusp, corner, a vertical tangent, or a discontinuity. y=3-cuberoot(x)

OpenStudy (anonymous):

the derivative of \[\sqrt[3]{x}\] is \[\frac{1}{3\sqrt[3]{x^2}}\] this has a vertical tangent at x = 0